# Evaluating Privacy Perceptions, Experience, and Behavior of Software Development Teams

### v1

@@ 6.3.2 Creation of Privacy Policies
Privacy policies describe how, why, and how long an application uses personal information. Regulations [18, 24] and the FTC [70] require companies to provide users with detailed privacy policies. Research shows that these policies may include inconsistencies with apps [76, 55], since they are either created by outside legal experts who may not fully comprehend the apps or by using privacy policy generators [75].
We asked the AD team about their experience and challenges with privacy policies (as other roles are often only indirectly involved). Out of the 70 participants, 3 did not provide any answer. Of the rest, only 11 (17%) have been involved in the creation of a privacy policy, and they used ‘legal experts’ the most (64.0%), followed by ‘templates’ (45.5%), and ‘privacy policy generators’ (36.4%). More than one response could be selected. Two of them mentioned that they either ‘search the Internet’ or ‘ask for team input’, in addition to using privacy policy generators and templates. In 60% (out of 45.5%) of cases that used ‘templates’, and in 50% (out of 36.4%) of cases that used ‘privacy policy generators’, a ‘legal expert’ has also been selected. This result matches with prior research that legal experts in a company are mainly involved in the privacy policy creation, which may lead to inconsistencies between the app and the policy [55]. We also noticed that those who said “Yes” are mostly from companies with less than 100 employees ($\sim$64%) and with a CPO ($\sim$55%).
Finally, we asked the 11 participants who responded ’Yes’, “What challenges did you face when creating your privacy policy?”. We received 10 responses. Six of them describe the challenges regarding compliance with regulations in multiple jurisdictions internationally, and understanding legal jargon, rules, and standards. One specifically had concerns regarding compliance, since they use privacy policy generators: “…differences between different countries and their requirements since we are international.” Five respondents describe their main challenge as ensuring completeness (i.e., covering all personal information), soundness, and language of privacy policies. E.g.: “Whether the wording I chose was going to cover all the bases I needed it to and whether it was clear and easy to understand.” or “Which rules and text to inform users; best guidelines.” One of those five respondents was also concerned about which template to choose. Four others did not find the process challenging since they trusted the legal expert to help.
@@ 6.3.3 Privacy Practices to Protect Users’ Data
We tailored some of the privacy practice questions based on the role, specifically for ISec, SD, and QA teams. We asked ISec members “How do you ensure that data collected from users is used only for intended purposes?”. They discussed various approaches. Some participants (32.08%) emphasized the importance of documentation to ensure transparency and accountability, with one noting “the meticulous documentation of every step in the data usage process”. Encryption emerged as a common theme, with participants mentioning sending encrypted documents and ensuring data is stored securely. A respondent states “I would send documents encrypted and compressed into a zip file, and instruct them to delete the file once the information is accessed.
Limiting access to data is another frequent approach, with 30.2% stressing the importance of restricting data access to only those who need it and maintaining logs to track any access. 16.98% stated the significance of transparency, ensuring they only collect necessary data, obtaining user consent, and regularly monitoring data usage. A few (9.43%) pointed out the importance of adhering to specific regulations, such as the Health Insurance Portability and Accountability Act (HIPAA) [28] and the Family Educational Rights and Privacy Act (FERPA) [44]. A few (16.98%) admitted to not having direct control over data but trusted their organization’s protocols and training to handle data responsibly.
@@
We asked the SD members “If you encounter a privacy concern at any point in the software development process, what steps would you take?”). More than 95% of them take the concerns very seriously. For example, one participant mentions “run a risk assessment” and another mentions “We take the app offline and start iteratively testing parts of the app to see where the privacy concern is.” About 36% deal with the concern internally to fix it and communicate it with the client and upper management. Another 35% directly escalate it to their supervisors, while 20% seek help from the ISec team or lawyers. A handful contact the client first.
We asked the QA team about third-party systems and their privacy compliance (i.e., “How do you verify that third-party systems used in your products are privacy compliant?”). Similarly, we received diverse responses. Only 56.60% confirmed that their companies conduct security audits of these third-party systems. Some mentioned the significance of conducting vulnerability assessments and penetration testing to ensure third-party systems’ compliance (23%). This is an important approach that ensures external software does not introduce vulnerabilities into the organization’s ecosystem that could lead to a potential breach of privacy if not audited. Some respondents discussed that they rely on reading privacy policies and contracts of third-party systems (28%), while others emphasized the importance of legal agreements and monitoring data transfers (15%). About 22% admitted to not being directly involved in this process, placing trust in their organization’s legal and security teams.
@@ 6.4 Privacy Awareness and Behaviors
We assess privacy behaviors w.r.t familiarity with regulations, PbD, PETs, and sources of privacy knowledge (i.e., RQ4).
@@ 6.4.1 Familiarity with Privacy Regulations
We asked participants “How did you learn about the previous regulations?”. More than one option could be selected. As shown in Table 4, a majority are self-taught while university education ranks second. Among all roles, the ISec team has the highest percentage of learning about regulations through university education (33.3%), which is more likely through cybersecurity courses. We also asked the participants to describe the other sources they used for learning about privacy regulations. In most cases, they mentioned ‘training at work’ as the source; however, 2 participants mentioned ‘social media’ and ‘YouTube’ as their source. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] GDPR is the most familiar regulation among all participants due to its comprehensiveness. ISec teams are more likely to learn about regulations through university education; hence, are more familiar with them than other groups. QA teams are the least familiar.
@@ 6.4.2 Familiarity with Privacy by Design (PbD)
Privacy by design (PbD) strategies introduced by Hoepman et al. [30] have gained interest in helping developers to be compliant with regulations. We asked the SD members (i.e., 198 participants) if they are aware of PbD, and if they answered yes, we asked them whether they used them in their work (see Appendix F - Table 17) and to list the ones they used. In general, less than half (46%) are familiar with PbD approaches while $\sim$25are unsure, which indicates the potential knowledge gap and opportunity for educating developers. Out of those who answered ‘yes’ to the awareness of the PbD question, only 57.1% had employed such strategies in their work. Of the remaining, 23.1% did not use them and 16.5% were unsure. This result suggests that even among developers who are familiar with such strategies, not everyone acts on this awareness – which may indicate the lack of usability and readiness of PbD for day-to-day developers’ tasks.
Lastly, we evaluated the responses about the usage of specific PbD strategies (multiple answers were possible). Interestingly, our results are aligned with the findings of Tahaei et al. [64] (see Table 5). Our top categories are ‘hide’ (22), ‘minimize’ (21), ‘inform’ (17), and ‘control’ (12), while ‘enforce’ (1) and ‘abstract’ (2) are rarely discussed. One participant mentions “Mostly minimise. Its the most straightforward.” This response reinforces our result in that ‘minimize’ is one of the easiest strategies to implement. We also received responses regarding Anne Cavoukian’s PbD principles [11] such as ‘privacy by default’ (6 times) and ‘proactive’ (twice). The use of PIA was also mentioned 5 times as a strategy. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Our findings show that PbD approaches are not yet commonly used and their lack of adoption underscores the gap in developers’ knowledge regarding PbD and their usability in day-to-day developers’ tasks.
@@ 6.4.3 Use of Privacy-Enhancing Technologies (PETs)
Using PETs is another critical component of privacy protection, that allows better protection and maintenance of data privacy against outside threats. We asked the SD team, who are the main users of PETs, if they used any PETs, and if so to list them. Out of the 198 participants, 2 did not respond. 111 of them (56.63%) mentioned that they use some PETs while 36 (18.37%) do not. About 25% are unsure. These results are almost aligned with the degree of familiarity and usage of PbD. There was an increase ($\sim$10%) in PETs familiarity and/or usage in comparison to PbD, which shows that these technologies are more common and tangible for developers, especially those related to encryption and access control. We grouped responses into 8 categories shown in Table 6 (definitions in Appendix G). Encryption and access control, which are primarily security-focused, were the most common, followed by anonymization methods and differential privacy.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] PETs are slightly more commonly used than PbD strategies. However, there is still a gap in their familiarity where more than 40% of developers do not use them or are unsure of their usage. The most commonly used PETs are more security-oriented concepts, than privacy.
Table 6: Usage of PETs in Software Development Process
Privacy Enhancing Technology (PET) Percentage
Data Protection and Encryption 70.48%
Access Control 34.29%
Anonymity and Pseudonymity 9.52%
Differential Privacy Approaches 8.57%
Secure Communication 7.62%
Privacy-Enhanced Web Browsing 1.90%
Privacy-Focused Hardware 0.95%
Privacy Governance & Compliance 0.95%
@@ 6.4.4 Developers’ Sources for Privacy Information
As discussed in Section 2, developers sometimes seek privacy-related guidance on forums, such as Reddit or Stack Overflow (SO). We asked the SD teams how often they use various developers’ forums for their privacy-related questions. Table 7 shows the distribution of the responses and their frequencies. $\sim$70and $\sim$58of the respondents use SO and GitHub at least 1-3 times per month, while for Reddit and Quora, this number is about 34.5% and 18.5%. About 57% of the respondents find these forums very or extremely useful, while less than 6% find them not useful at all. In cases where they do not find the answer on these forums, the SD team discusses their questions with the security or privacy experts, asks their teammates, or uses Artificial intelligence tools. In Appendix I, we provide a more detailed analysis regarding the usage of the forums with respect to various factors.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Developers often seek privacy-related information from online forums, where more than 50% of participants use either Stack Overflow or GitHub 1-3 times per month and they find these forums useful.
@@ 7 Location Analysis
Our large-scale survey included two key regions: the US (189 responses), and non-US (173 responses from 22 countries: EU+UK, South Africa, Canada+Mexico (C+M), and South America), to examine differences in perceptions, experiences, and behaviors. To evaluate the difference in perception, we examine whether participants’ location (US or non-US) correlates with their confidence in privacy and security measures (H6a) and the presence of a CPO (H6b). As shown in Appendix J, both hypotheses do not hold ($p-values$ are 0.1896 and 0.2727). Table 8 shows the presence of a CPO based on the location. The number of ‘Yes’ is almost equal between the US and EU+UK while slightly more US participants mentioned “no CPO” than those in EU+UK. This is not surprising since GDPR, the Data Protection Act of 2018 from the EU, and US HIPAA (Art.164.530) all have a requirement for having a privacy officer or officials in similar role. Interestingly, for South Africa (SA), $\sim$70of the participants said ‘Yes’. The SA Protection of Personal Information Act (POPIA) [1] - Chapter 5 enforces having an information officer.
@@
South Africa 69.2% 23.1% 7.7% 0%
C+M 20% 40% 40% 0%
South America 0% 50% 50% 0%
We evaluated whether there is a significant difference between participants’ experience in the two regions regarding the creation of PIA (H6c) and the number of privacy breaches (H6d). With $p-values$ 0.5870 and 0.2016, we find no correlation. We also analyzed the privacy behaviors in the two regions w.r.t familiarity with PbD (H6e) and usage of PETs (H6f). With $p-values$ 0.6246 and 0.2637, we do not find any correlation that suggests that usage of PETs and PbD are equally (un)common in both regions.
Since participants are from regions governed by different privacy laws, we investigated their familiarity with CCPA [24] (H6g) and GDPR [18] (H6h). As expected, we find a significant correlation between the participant’s familiarity with the two regulations (CCPA $p-value<0.0001$, GDPR $p-value<0.0001$). The results for these eight tests do not differ if we compare the US with only the EU and UK.
Due to the global reach of many apps, SDLC teams are responsible for complying with various regulations. We further evaluated the responses to the familiarity with each regulation between the US vs non-US. We combined the responses given for at least somewhat familiarity (i.e., somewhat, very, extremely familiar) and found that participants in the US and South Africa are most familiar with HIPAA while the rest are most familiar with GDPR. Those from US and South Africa are also more familiar with CCPA and CPRA than others. Table 9 shows the distribution. South American participants were not very familiar with any of the regulations.
@@ 8 Discussion
Summary of Findings With respect to privacy perception, our survey identifies that the majority of the participants define privacy in terms of control over personal information and disclose only when needed, or in terms of security. In other research [32, 62] data protection and security were the most common definitions. Having a CPO or a similar role positively impacts confidence in protecting users’ data. However, we found out that a sizable portion of the participants are unaware of such a role in their company, which may lead to ineffectiveness in utilizing privacy tools or reducing privacy breaches. Lack of proper communication among various roles is a challenge that other research also identified [32, 62]. Our findings also align with Balebako et al.’s 2014 study [6] and Horstmann et al. [32], which observed a correlation between company size and having a CPO. However, we did not observe significant location-based differences in these perceptions. This is interesting but not surprising since GDPR, HIPAA, the UK Data Protection Act, and POPIA all require a CPO or similar roles. Several of our US participants mentioned (in Q27) that they collect Protected Health Information (PHI), which falls under HIPAA; e.g., one participant says “Health related data about people involved with our insurance companies”. The extensive privacy requirements from these regulations likely explain why we observed no significant geographical differences in terms of participants’ confidence, familiarity with PbD, and the usage of PETS.
@@
In terms of privacy experience, most participants rely on legal experts to help create privacy policies; unlike [6] where creating a privacy policy was not the priority. Our study also shows that participants are primarily concerned about multi-jurisdictional compliance. Most of them are not involved in creating a PIA. The majority of those involved believe a PIA should be created during the planning or analysis phases, but some still create a PIA during development, testing, and deployment; this is almost similar to findings in [26, 32]. Our participants emphasized the importance of detailed documentation regarding the data lifecycle, as well as using encryption and access control tools to protect the confidentiality and integrity of data. Interestingly, the QA teams rely more than others on security, privacy, and legal experts to implement and enforce privacy and security rules. Other studies did not examine the privacy practices of QA roles, separately.
@@
Regarding privacy behavior, we identified that less than half of the participants are aware of PbD and an even smaller number use them. Similar to[64], ‘hide’, ‘minimize’, ‘inform’, and ‘control’ are more commonly used. The usage of PETs is slightly more prevalent than PbD, but the focus is more on security practices, such as encryption and access control; similar to other research that found security concepts are more tangible [26, 6, 62, 32]. Anonymization techniques are not used frequently enough. We also find that although $\sim 52\%$ of our participants are from the US, most are more familiar with GDPR than US-based regulations such as COPPA and CCPA. ISec experts are among the most knowledgeable about various regulations, while QA teams are the least familiar. Other work focus mainly on GDPR and CCPA and do not explore details regarding participants’ familiarity [26, 6, 62, 32]. Most participants tend to seek answers to their privacy questions from developers’ forums in addition to legal/policy experts; unlike [6] where they used ‘friends’ or ‘social media’.
@@
Research Directions Insights from the related work and our survey results highlight the need for approaches to operationalize PbD strategies and incorporate them into design and development. PbD patterns [71, 72] provide detailed information about their usage and high-level solutions but still lack implementation. Approaches that detect privacy behaviors in code [37, 35] and further link them to patterns or leverage automated code generation techniques to generate code from privacy patterns are yet to be explored.
@@
Our survey highlights software teams’ challenges while creating PIAs and privacy policies. Research directions that focus on automated approaches to detect the information types, privacy practices, and purposes pre- [33] and post-development [37, 35], or to generate privacy statements directly from code [36] could aid software teams in creating more accurate PIA and privacy notices – which alleviate the challenges regarding consistency and compliance.
@@
Our survey indicates that software teams face challenges in understanding and adhering to privacy regulations; thus, there is a need for approaches to help better understand such regulations, and establish and maintain compliance. However, most research focuses on detailed requirements analysis, not suitable for agile app development. Future studies could focus not only on automated extraction of legal/privacy requirements but also on generating (privacy-related) user stories to be used in agile development. Research directions on automated approaches to monitor compliance and nudge developers towards compliant approaches are also worth addressing.
Educational Takeaway Similar to other work [6, 32, 62], our work shows the need for a more focused educational approach toward privacy in the SDLC. While currently, many courses emphasize security, it is important to tailor specific courses that include advanced privacy topics such as: regulations; the importance of PIA and other artifacts; challenges in notice, choices, and privacy policy creation; and approaches such as PbD, differential privacy, and federated learning. This distinction between privacy from security is crucial since privacy encompasses a broad spectrum of concerns, including data handling, user consent, and transparency.
Software teams should be equipped with educational modules and tool supports that foster and support life-long learning of dynamic privacy concepts. Nudging developers towards more privacy-preserving solutions through online support and tools is important. Balebako et al. [5] suggest that with the right guidance, developers can be encouraged to prioritize privacy in their design and development processes.

### v2

@@ 6.3.2 Creation of Privacy Policies
Privacy policies describe how, why, and how long an application uses personal information. Regulations [27, 35] and the FTC [92] require companies to provide users with detailed privacy policies. Research shows that these policies may be inconsistent with apps [99, 78], since they are either created by outside legal experts (who may not fully comprehend the apps) or by using privacy policy generators [98].
We asked the AD team about their experience and challenges with privacy policies (as other roles are often only indirectly involved). Out of the 70 participants, 3 did not provide any answer. Of the rest, only 11 (17%) have been involved in the creation of a privacy policy, and they used ‘legal experts’ the most (64.0%), followed by ‘templates’ (45.5%), and ‘privacy policy generators’ (36.4%). More than one response could be selected. Two of them mentioned that they either ‘search the Internet’ or ‘ask for team input’, in addition to using privacy policy generators and templates. In 60% (out of 45.5%) of cases that used ‘templates’, and in 50% (out of 36.4%) of cases that used ‘privacy policy generators’, a ‘legal expert’ has also been selected. This result matches with prior research that legal experts in a company are mainly involved in the privacy policy creation, which may lead to inconsistencies between the app and the policy [78]. We also noticed that those who said ‘Yes’ are mostly from companies with less than 100 employees ($\sim$64%) and with a CPO ($\sim$55%).
Finally, we asked the 11 participants who responded ‘Yes’, “What challenges did you face when creating your privacy policy?”. We received 10 responses. Six of them describe the challenges regarding compliance with regulations in multiple international jurisdictions, and understanding legal jargon, rules, and standards. One specifically had concerns regarding compliance, since they use privacy policy generators: “…differences between different countries and their requirements since we are international.” Five respondents describe their main challenge as ensuring completeness (i.e., covering all personal information), soundness, and language of privacy policies. E.g.: “Whether the wording I chose was going to cover all the bases I needed it to and whether it was clear and easy to understand.” or “Which rules and text to inform users;…” One of those five respondents was also concerned about which template to choose. Four others did not find the process challenging since they trusted the legal expert to help.
@@ 6.3.3 Privacy Practices to Protect Users’ Data
We tailored some of the privacy practice questions based on the role, specifically for ISec, SD, and QA teams. We asked ISec members: “How do you ensure that data collected from users is used only for intended purposes?”. They discussed various approaches. $\sim$32emphasized the importance of documentation to ensure transparency and accountability, with one noting “the meticulous documentation of every step in the data usage process”. Encryption emerged as a common theme, with participants mentioning sending encrypted documents and ensuring data is stored securely. A respondent states “I would send documents encrypted and compressed into a zip file, and instruct them to delete the file once the information is accessed.” Limiting access to data is another frequent approach, with 30.2% stressing the importance of restricting data access to only those who need it and maintaining logs to track any access. 16.98% stated the significance of transparency, ensuring they only collect necessary data, obtaining user consent, and regularly monitoring data usage. A few (9.43%) pointed out the importance of adhering to specific regulations, such as the Health Insurance Portability and Accountability Act (HIPAA) [43] and the Family Educational Rights and Privacy Act (FERPA) [64]. 16.98% admitted to not having direct control over data but trusted their organization’s protocols and training to handle data responsibly.
@@
We asked the SD members: “If you encounter a privacy concern at any point in the software development process, what steps would you take?”. More than 95% of them take the concerns very seriously. For example, one participant mentions “run a risk assessment” and another mentions “We take the app offline and start iteratively testing parts of the app to see where the privacy concern is.” About 36% deal with the concern internally to fix it and communicate it with the client and upper management. Another 35% directly escalate it to their supervisors, while 20% seek help from the ISec team or lawyers. A handful contact the client first.
We asked the QA team: “How do you verify that third-party systems used in your products are privacy compliant?”. Similarly, we received diverse responses. Only 56.6% confirmed that their companies conduct security audits of these third-party systems. Some mentioned the significance of conducting vulnerability assessments and penetration testing to ensure third-party systems’ compliance (23%). Some respondents discussed that they rely on reading privacy policies and contracts of third-party systems (28%), while others emphasized the importance of legal agreements and monitoring data transfers (15%). About 22% admitted to not being directly involved in this process, placing trust in their organization’s legal and security teams, which is aligned with findings in [19].
@@ 6.4 Privacy Awareness and Behaviors
We assess privacy behaviors based on familiarity with regulations, PbD, PETs, and such knowledge sources (i.e., RQ4).
@@ 6.4.1 Familiarity with Privacy Regulations
We asked participants “How did you learn about the previous regulations?”. More than one option could be selected. As shown in Table 4, the majority are self-taught while university education ranks second. Among all roles, the ISec team has the highest percentage of learning about regulations through university education (33.3%), which is more likely through cybersecurity courses. We also asked the participants to describe the other sources they used to learn about privacy regulations. In most cases, they mentioned ‘training at work’ as the source; however, 2 participants mentioned ‘social media’ and ‘YouTube’ as their source. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] GDPR is the most familiar regulation among all participants due to its comprehensiveness. ISec teams are more likely to learn about regulations through university education; hence, are more familiar with them than other groups. QA teams are the least familiar.
@@ 6.4.2 Familiarity with Privacy by Design (PbD)
Privacy by design (PbD) strategies introduced by Hoepman et al. [45] have gained interest in helping developers to be compliant with regulations. We asked the SD members (i.e., 198 participants) if they are aware of PbD, and if they answered yes, whether they used them (see Appendix G - Table 17) and to list the ones they used. $\sim$46are familiar with PbD approaches while $\sim$25are unsure, which indicates the potential knowledge gap and opportunity for educating developers. Out of those who answered ‘Yes’ to the awareness of the PbD question, only 57.1% had employed such strategies in their work. Of the remaining, 23.1% did not use them and 16.5% were unsure. This result suggests that even among developers who are familiar with such strategies, not everyone acts on this awareness – which may indicate the lack of usability and readiness of PbD for day-to-day developers’ tasks [88] or other organizational factors, such as lack of resources [51].
Lastly, we evaluated the responses about the usage of specific PbD strategies (multiple answers were possible). Interestingly, our results are aligned with the findings of Tahaei et al. [87] (see Table 5). Our top categories are ‘hide’ (22), ‘minimize’ (21), ‘inform’ (17), and ‘control’ (12), while ‘enforce’ (1) and ‘abstract’ (2) are rarely discussed. One participant mentions “Mostly minimise. Its the most straightforward.” This response reinforces our result in that ‘minimize’ is one of the easiest strategies to implement. We also received responses regarding Anne Cavoukian’s PbD principles [17] such as ‘privacy by default’ (6 times) and ‘proactive’ (twice). The use of PIA was also mentioned 5 times as a strategy. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Our findings show that PbD approaches are not yet commonly used, and their lack of adoption underscores the gap in developers’ knowledge regarding PbD and their usability in day-to-day developers’ tasks.
Table 6: Usage of PETs in Software Development Process
Privacy Enhancing Technology (PET) Percentage
Encryption 70.48%
Access Control/Identity Protection 34.29%
Anonymity and Pseudonymity 9.52%
Differential Privacy Approaches 8.57%
Secure Communication/VPN 8.57%
Privacy-Enhanced Anti Web Tracking 0.0%
@@ 6.4.3 Use of Privacy-Enhancing Technologies (PETs)
Using PETs is another critical component of privacy protection, that allows better protection and maintenance of data privacy against outside threats. We asked the SD team, who are the main users of PETs, if they used any PETs, and if so to list them. Out of the 198 participants, 2 did not respond. 111 of them (56.63%) mentioned that they use some PETs while 36 (18.37%) do not. About 25% are unsure. These results are almost aligned with the degree of familiarity and usage of PbD. There was an increase ($\sim$10%) in PETs familiarity and/or usage in comparison to PbD, which shows that these technologies are more common and tangible for developers, especially those related to encryption and access control. We grouped responses into 6 categories shown in Table 6 (definitions in Appendix K). Encryption and access control, which are primarily security-focused, were the most common, followed by anonymization methods and differential privacy.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] PETs are slightly more commonly used than PbD strategies. However, there is still a gap in their familiarity, where more than 40% of developers do not use them or are unsure of their usage. The most commonly used PETs are more security-oriented concepts, than privacy.
@@ 6.4.4 Developers’ Sources for Privacy Information
As discussed in Section 2, developers sometimes seek privacy-related guidance on forums, such as Reddit or Stack Overflow (SO). We asked the SD teams how often they use various developers’ forums for their privacy-related questions. Table 7 shows the distribution of the responses and their frequencies. $\sim$70and $\sim$58of the respondents use SO and GitHub at least 1-3 times per month, while for Reddit and Quora, this number is about 34.5% and 18.5%. About 57% of the respondents find these forums very or extremely useful, while less than 6% find them not useful at all. In cases where they do not find the answer on these forums, the SD team discusses their questions with the security or privacy experts, asks their teammates, or uses AI tools. In Appendix I, we provide a more detailed analysis regarding the usage of the forums.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Developers often seek privacy-related information from online forums, where more than 50% of participants use either Stack Overflow or GitHub at least 1-3 times per month and they find these forums useful.
@@ 7 Location Analysis
Our large-scale survey has responses from the US (189 responses) and non-US (173 responses from 22 countries: EU+UK, South Africa, Canada (CA), Mexico, and Chile), enabling us to examine differences in perceptions, experiences, and behaviors. We group the countries into three regions based on their similarities in privacy regulations: US+CA (192), EU+UK (132), and ‘Other’ countries (38). To evaluate the difference in perception, we examine whether participants’ location correlates with their confidence in privacy and security measures (H6a in Appendix J) and the presence of a CPO (H6b). Both hypotheses do not hold ($p-values$ are 0.0567 and 0.6470). Table 8 shows the presence of a CPO across the three regions. The percentage of ‘Yes’ is almost equal between US+CA, EU+UK, and the ‘Other’ countries, while slightly more US+CA participants mentioned “no CPO” than elsewhere. This is not surprising since GDPR, the UK Data Protection Act of 2018, and the US HIPAA (Art.164.530) all require having a privacy officer or officials in a similar role.
@@
Other Countries 43.5% 30.4% 26.1% 0%
We evaluated whether there is a significant difference between participants’ experience in the three regions regarding the creation of PIA (H6c) and the number of privacy breaches (H6d). With $p-value$ 0.7724, we find no correlation for PIA. However, there is a correlation between the regions and privacy breaches ($p-value=$ 0.0010). We also analyzed the privacy behaviors in the three regions concerning familiarity with PbD (H6e) and usage of PETs (H6f). With $p-values$ 0.3120 and 0.8588, we do not find any correlation that suggests that usage of PETs and PbD are equally (un)common in all regions. Since participants are from regions governed by different privacy laws, we investigated their familiarity with CCPA [35] (H6g) and GDPR [27] (H6h). As expected, we find a significant correlation between the participant’s familiarity with the two regulations ($p-values$ are $<0.0001$ and $0.0009$ respectively). Due to the global reach of many apps, SDLC teams are responsible for complying with various regulations. We further evaluated the responses to the familiarity with each regulation in various regions. We combined the responses given for at least somewhat familiarity (i.e., somewhat, very, extremely familiar) and found that participants in the US+CA are most familiar with HIPAA while the rest are most familiar with GDPR. Those from ‘Other’ countries are also more familiar with the US regulations than those residing in the EU+UK. Table 9 shows the distribution.
@@ 8 Discussion
Summary of Findings Concerning privacy perception, our survey identifies that the majority of the participants define privacy in terms of control over personal information and disclose only when needed, or in terms of security. In other research [48, 85], data protection and security were the most common definitions. Having a CPO or a similar role positively impacts confidence in protecting users’ data. However, we found out that a sizable portion of the participants are unaware of such a role in their company, which may lead to ineffectiveness in utilizing privacy tools or reducing privacy breaches. Lack of proper communication among various roles is a challenge that other research also identified [48, 85]. Our findings also align with [7] and [48], which observed a correlation between company size and having a CPO. However, we did not observe significant location-based differences in these perceptions. This is interesting but not surprising, since GDPR, HIPAA, the UK Data Protection Act, and Protection of Personal Information Act (POPIA) all require a CPO or similar roles. Several of our US participants mentioned (in Q27) that they collect Protected Health Information (PHI), which falls under HIPAA; e.g., one participant says “Health related data about people involved with our insurance companies”. The extensive privacy requirements from these regulations likely explain why we observed no significant geographical differences in terms of participants’ confidence, familiarity with PbD, and the usage of PETS.
@@
In terms of privacy experience, most participants rely on legal experts to help create privacy policies; unlike [7] where creating a privacy policy was not the priority. Our study also shows that participants are primarily concerned about multi-jurisdictional compliance. Most of them are not involved in creating a PIA. The majority of those involved believe a PIA should be created during the planning or analysis phases; this is almost similar to findings in [40, 48]. Our participants emphasized the importance of detailed documentation regarding the data lifecycle, as well as using encryption and access control tools to protect the confidentiality and integrity of data. Interestingly, the QA teams rely more than others on security, privacy, and legal experts to implement and enforce privacy and security rules. Other studies did not examine the privacy practices of QA roles, separately.
@@
Regarding privacy behavior, we identified that less than half of the participants are aware of PbD and an even smaller number use them. Similar to[87], ‘hide’, ‘minimize’, ‘inform’, and ‘control’ are more commonly used. The usage of PETs is slightly more prevalent than PbD, but the focus is more on security practices, such as encryption and access control; similar to other research that found security concepts are more tangible [40, 7, 85, 48]. Anonymization techniques are not used frequently enough. We also find that although $\sim 53\%$ of our participants are from the US+CA, most are more familiar with GDPR than US-based regulations such as COPPA and CCPA. ISec experts are among the most knowledgeable about various regulations, while QA teams are the least familiar. Other works focus mainly on GDPR and CCPA and do not explore details regarding participants’ familiarity [40, 7, 85, 48]. Most participants tend to seek answers to their privacy questions from developers’ forums in addition to legal/policy experts; unlike [7] where they used ‘friends’ or ‘social media’.
@@
Research Directions Insights from the related work and our survey results highlight the need for approaches to operationalize PbD strategies and incorporate them into design and development. PbD patterns [94, 93] provide detailed information about their usage and high-level solutions, but still lack implementation. Approaches that detect privacy behaviors in code [55, 53] and further link them to patterns, or leverage automated code generation techniques to generate code from privacy patterns are yet to be explored.
@@
Our survey highlights software teams’ challenges in creating accurate PIAs and privacy policies. Research directions that focus on automated approaches to detect the information types, privacy practices, and purposes pre- [49] and post-development [55, 53], or to generate privacy statements from code [54] could alleviate the challenges regarding accuracy, consistency, and compliance.
@@
Our survey indicates that software teams face challenges in understanding and adhering to privacy regulations; thus, there is a need for approaches to help better understand such regulations, and establish and maintain compliance. However, most research focuses on detailed requirements analysis, not suitable for agile app development. Future studies could focus not only on automated extraction of legal/privacy requirements but also on generating (privacy-related) user stories to be used in agile development. Research directions on automated approaches to monitor compliance and nudge developers towards compliant approaches are also worth addressing [18].
Educational Takeaway Similar to other work [7, 48, 85], our work shows the need for a more focused educational approach toward privacy in the SDLC. While currently, many courses emphasize security, it is important to tailor specific courses that include advanced privacy topics such as: regulations; the importance of PIA and other artifacts; challenges in privacy policy creation; and approaches such as PbD, differential privacy, and federated learning. This distinction between privacy from security is crucial since privacy encompasses a broad spectrum of concerns, including data handling, user consent, and transparency. Software teams should be equipped with educational modules and tools that foster and support life-long learning of dynamic privacy concepts. Nudging developers towards more privacy-preserving solutions through online support and tools is important. Balebako et al. [6] suggest that with the right guidance, developers can be encouraged to prioritize privacy in their design and development processes.